# Spline-FRIDA: Towards Diverse, Humanlike Robot Painting Styles with a Sample-Efficient, Differentiable Brush Stroke Model

### v1

@@ VI-C3 Traj2Stroke
Our new method described in Section V-D.
@@ VI-C4 Traj2Stroke with U-Net
Our new method, but with an additional Pix2Pix network appended after the output layer. The goal of this additional layer is to refine the Traj2Stroke output by learning subtle effects such as texture and bristle drag. The architecture of the new layer closely follows that of U-Net [26]. We freeze the U-Net weights during the first half of the training and unfreeze them for the second half. The purpose of this is to train the Traj2Stroke portion first and get it as close as possible to the ground truth, before using the U-Net to refine it. Inspired by the success of ControlNet [27], we utilize a zero-convolution so that the U-Net has no effect on the output while its weights are frozen.
An important metric we wish to evaluate is generalizability. We define a stroke model as generalizable if it can output reliable stroke predictions, even when given out-of-training-distribution trajectories. This is important because stroke models are time-consuming to train; it requires physically executing trajectories on a robot. A generalizable stroke model only needs to be trained once per drawing medium. It can then be used out-of-the-box for planning with any TrajVAE, thus saving time and effort.
In order to evaluate generalizability, we train and test the stroke model on trajectories from different distributions. More precisely, we create two datasets, A and B. Both datasets contain (trajectory, stroke image) pairs. For dataset A, the trajectories are sampled from a generic TrajVAE, trained on a session that we judge to have good stroke diversity. For dataset B, we use trajectories from more specialized TrajVAEs, trained on sessions with very unique styles. We train the model using dataset A, and we evaluate generalizability by checking its performance on dataset B.
@@
Visually, example predictions generated by each model can be seen in Figure 8. All examples are from dataset B, meaning that these trajectories are from a TrajVAE different from the one used during training. The vanilla CNN fails to generalize in certain cases. Changing standard convolution layers into CoordConv layers does not seem to help much. The Traj2Stroke model is near-perfect for the sharpie strokes. It also generates reasonable output for the brush strokes, although it is unable to capture the texture. The U-Net layer is an attempt to mitigate this; its design aims to combine the precision of Traj2Stroke with the flexibility of convolutions. While the outputs indeed seem to capture the texture of paint, they also tend to be more blotchy, which is concerning for an optimization-based planner.
@@ VII Limitations/Future Work
Artistic style is nuanced and subjective. Style can be broken down into roughly 7 elements: Lines, shapes, space, texture, value, and color [2]. In this paper, we focus on capturing style by modeling brush stroke trajectories. These are intra-stroke attributes which only cover the style elements of lines and shapes. In future work, we hope to adapt the method to capture the other, inter-stroke style elements.
Spline-FRIDA is computationally expensive, e.g., 400-stroke paintings can take approximately one hour to optimize plans for on a machine with an NVIDIA GeForce 4090. To adhere to hardware memory constraints, we render only a subset of strokes (we use $n=80$ in practice) while recording gradients with each forward pass. Beyond computational inneficiencies, the optimization methods for stroke-based rendering in general are susceptible to local minima. Results are dependent on stroke parameter initialization. In future work, we would like to explore training models, such as reinforcement learning agents, to perform stroke prediction. This could improve runtime planning speeds, memory constraints, and potentially help avoid local minima.

### v2

@@ VI-C3 Traj2Stroke
This is our main method, with the rule-based transformations, that was described in Section V-D.
@@ VI-C4 Traj2Stroke with U-Net
Our main method, but with an additional convolutional network attached after the output layer. The goal of this additional network is to refine the Traj2Stroke output by learning subtle effects such as texture and bristle drag. Its architecture closely follows that of U-Net [28]. We freeze the U-Net weights during the first half of the training and unfreeze them for the second half. The purpose of this is to train the Traj2Stroke portion first and get it as close as possible to the ground truth, before using the U-Net to refine it. Inspired by the success of ControlNet [29], the U-net is initialized with a zero-convolution final transformation so that it initially performs the identity function.
Since generating training strokes and training a new dynamics model for every new stroke style is time consuming, the ability of the dynamics model to generalize to unseen styles is important. In order to evaluate generalizability, we train and test the stroke model on trajectories from different distributions. More precisely, we create two datasets, A and B. Both datasets contain (trajectory, stroke image) pairs. For dataset A, the trajectories are sampled from a generic TrajVAE, trained on a session that we judge to have good stroke diversity. For dataset B, we use trajectories from more specialized TrajVAEs, trained on sessions with very unique styles. We train the model using dataset A, and we evaluate generalizability by checking its performance on dataset B.
@@
Visually, example predictions generated by each model can be seen in Figure 9. All examples are from dataset B, meaning that these trajectories are out of distribution from the training set. The vanilla CNN with and without CoordConv fails to generalize in certain cases. The Traj2Stroke model performs near-perfect for the sharpie strokes and captures the general shape of the brush strokes. Adding the U-Net to Traj2Stroke helped capture the texture of the brush strokes, but the added parameters hurt generalization to very out-of-distribution strokes, such as the star example.
@@ VII Limitations/Future Work
In this paper, we focus on capturing style by modeling brush stroke trajectories. However, these are intra-stroke attributes and we do not consider other aspects of style, including inter-stroke style elements.
Additionally, Spline-FRIDA is computationally expensive (400-stroke paintings take approximately one hour to plan for on a machine with an NVIDIA GeForce 4090). Hardware memory constraints also force us to optimize only a subset (80 in practice) of strokes at a time. Beyond computational inneficiencies, the optimization methods for stroke-based rendering in general are susceptible to local minima. Results are dependent on stroke parameter initialization. In future work, we would like to explore training models, such as reinforcement learning agents, to perform stroke prediction. This could improve runtime planning speeds, memory constraints, and potentially help avoid local minima.
@@ VIII Conclusions
IX Acknowledgements
We thank Howard Halim for his early ideas related to the Traj2Stroke model. This work was supported in part by the Technology Innovation Program (20018295, Meta-human: a virtual cooperation platform for specialized industrial services) and funded by the Ministry of Trade, Industry & Energy (MOTIE, Korea) and NSF IIS-2112633.